Some recent financial statements for Smolira Golf Corp. follow. SMOLIRA GOLF CORP. 2014 and 2015 Balance Sheets Assets Liabilities and Owners’ Equity 2014 2015 2014 2015 Current assets Current liabilities Cash $ 23,066 $ 25,300 Accounts payable $ 24,384 $ 28,300 Accounts receivable 13,648 16,400 Notes payable 14,000 12,000 Inventory 27,152 28,300 Other 12,771 19,700 Total $ 63,866 $ 70,000 Total $ 51,155 $ 60,000 Long-term debt $ 81,000 $ 90,000 Owners’ equity Common stock and paid-in surplus $ 55,000 $ 55,000 Accumulated retained earnings 213,406 229,000 Fixed assets Net plant and equipment $ 336,695 $ 364,000 Total $ 268,406 $ 284,000 Total assets $ 400,561 $ 434,000 Total liabilities and owners’ equity $ 400,561 $ 434,000 SMOLIRA GOLF CORP. 2015 Income Statement Sales $ 333,706 Cost of goods sold 215,200 Depreciation 49,300 Earnings before interest and taxes $ 69,206 Interest paid 15,500 Taxable income $ 53,706 Taxes (30%) 16,112 Net income $ 37,594 Dividends $ 22,000 Retained earnings 15,594 Smolira Golf Corp. has 40,000 shares of common stock outstanding, and the market price for a share of stock at the end of 2015 was $42. What is the price-earnings ratio? (Do not round intermediate calculations and round your final answer to 2 decimal places, e.g., 32.16.) Price-earnings ratio times What are the dividends per share? (Round your answer to 2 decimal places, e.g., 32.16.) Dividends per share What is the market-to-book ratio at the end of 2015? (Round your answer to 2 decimal places, e.g., 32.16.) Market-to-book ratio times If the company’s growth rate is 10 percent, what is the PEG ratio? (Do not round intermediate calculations and round your final answer to 2 decimal places, e.g., 32.16.) PEG ratio times
Answer #1

Price Earning Ratio : current market price of stock by its earning per share. It shows stock performance which tells how much money one needs to pay for each worth of earnings.

P-E ratio = 42 * 40000/37594 = 44.69

Dividend per share = it tells how much dividend company pay to its share holder. It derived by total dividend payout/no. of share

Dividend per share = 22,000/40000 = $.55

Market to book ratio : find the value of firm by comparing market value of firm to its book value. Low market to book ratio (less than 1) tells that stock is undervalued and vice versa.

market to book ratio = 42 * 40000/(55000 + 229000) = 5.92

PEG ratio: PE ratio by growth of company. It is used to determone company stock value while taking the company's earning growth in account.

PEG ratio = 44.69/10 = 4.47